smoore69:

How much of an alloy with 30% copper must be added to 25 pounds of a second alloy which is 20% copper to create a 27% copper alloy?

smoore69:

Alexis1415:

The answer is: 58.33 lb, There are already 25 lbs of alloy copper, of this 25 lbs 20% is pure copper content. 25(0.5)= 5 lbs

shortyashawty000056:

he is right

Alexis1415:

she*

shortyashawty000056:

sorry she

surjithayer:

let x pounds of 30 % copper ne added to 25 pounds of 20% copper to form (x+25) pounds of 27% copper. \[x \times 30 \%+25\times 20 \%=(x+25)\times 27 \% \] multiply by 100 \[30x+25\times20=27x+25\times27\] \[30 x-27 x=25\times 27-25\times 20=25\times (27-20)\] \[3 x=25\times 7=175\] \[x=\frac{ 175 }{ 3 }\approx 58.33 ~pounds\]